# Improving Handwritten Text Recognition via 3D Attention and Multi-Scale Training

### v2

@@ 3.2.2 Overall Comparison
Finally, Table 6 shows an overall comparison of our proposed method and other state-of-the-art methods on different test sets. Without using external data, our proposed network can make a new milestone. Compared with the best networks trained by using external data, the proposed network can also keep comparable results on all datasets. For the listed results on the IAM dataset, the standard partition and no word lexicons were used. Please note there exists an unofficial partition (RWTH partition) of the IAM dataset. It is widely used by many researchers, their results are not compared in this paper.
@@ 3.2.3 Visualization Analysis
Figure 6: The typical samples of the attention visualization on the SCUT-HCCDoc dataset.
Figure 7: The visualization of the 3D attention weights on the IAM images.
Figure 8: The typical samples of the attention visualization on the SCUT-EPT dataset.
4 Conclusion
In this paper, main canonical neural units including attention mechanisms, fully-connected layer, recurrent unit and convolutional layer are ingeniously organized into a network. The network weights are effectively optimized by the multi-scale training. Experimental results show that the proposed method can achieve comparable results with the state-of-the-art methods. Besides, the visualization analysis demonstrates the proposed attention mechanism is reasonable. For future work, we will investigate large vision and language models for document analysis.
@@ Acknowledgments
This work was supported by the National Natural Science Foundation of China under Grant No. 62106031. The author would like to thank Professor Jun Du of the University of Science and Technology of China (USTC). When the author was a Ph.D. candidate at USTC, we successfully applied the HMM and the deep neural network to the text recognition task. The supervisor (Prof. Jun Du) always said, can we combine three mainly segmentation-free methods? Although the HMM is not explicitly used in this paper, the question inspired the author to delve into this problem while self-isolated at home. The author also would like to thank Yan Wang for applying the code of extracting text lines from original SCUT-HCCDoc documents.

### v3

@@ 3.2.2 Overall Comparison
Finally, Table 6 shows an overall comparison of our proposed method and other state-of-the-art methods on different test sets. Without using external data, our proposed network can make a new milestone. Compared with the best networks trained by using external data, the proposed network can also achieve comparable results on all datasets. For the listed results on the IAM dataset, the standard partition and no word lexicons were used. Please note that there exists an unofficial partition (RWTH partition) of the IAM dataset. It is widely used by many researchers, their results are not compared in this paper.
@@ 3.2.3 Visualization Analysis
Figure 7: The typical samples of the attention visualization on the SCUT-HCCDoc images [31].
Figure 8: The visualization of the 3D attention weights on the IAM images [33].
Figure 9: The typical samples of the attention visualization on the SCUT-EPT images [32].
4 Discussion and Conclusion
From the above experiments, we can observe that the proposed network has achieved good results on multiple handwritten text datasets. One important factor is that the proposed network can absorb the advantages of three typical segmentation-free sequence recognition algorithms. Compared with the HMM, which relies on artificially designed features and the generative model (GMM-HMM) for character modeling, we do not need to perform complex character modeling and inference. However, we have integrated the state-level feature extraction method based on sliding windows in HMM into the intermediate process of the network, and ingeniously extracted multi-scale block features for the training of the network, which potentially increases the training data and enables the network to better handle characters of different sizes. Compared with the encoder-decoder method based on the attention mechanism, which is prone to attention drift when applied to long handwritten text recognition, the proposed 3D attention mechanism is conducted on the sliding window to generate better local visual features. Whether from intuition or from our experimental results, the proposed 3D attention can better take into account the 2D natural properties of characters. Without too many tricks, we use the most popular neural units to implement the 3D attention. The extracted sequence visual features are used to obtain global and local context features through the self-attention mechanism and the LSTMs, respectively. Obviously, the representation after combining visual features and contextual features is suitable for the text recognition task. Finally, by cascading CTC loss and CE loss, the whole network can be optimized well. As shown in Fig 2, in this paper, during the test stage, we only use one of the network branches and only use the faster CTC method, without the cyclic decoding way of common ED methods.
Overall, main canonical neural units including attention mechanisms, fully-connected layers, recurrent units and convolutional layers are ingeniously organized into a network. The network weights are effectively optimized by the multi-scale training. Experimental results show that the proposed method can achieve comparable results with the state-of-the-art methods. Besides, the visualization analysis demonstrates the proposed attention mechanism is reasonable. For future work, we will investigate large vision and language models for document analysis.
@@ Acknowledgments
This work was supported by the National Natural Science Foundation of China under Grant No. 62106031. The author would like to thank Professor Jun Du of the University of Science and Technology of China (USTC). When the author was a Ph.D. candidate at USTC (September 2015 to June 2020), we successfully applied the HMM and the deep neural network to the text recognition task. The supervisor (Prof. Jun Du) always said, can we combine three mainly segmentation-free methods? Although the HMM is not explicitly used in this paper, the question inspired the author to delve into this problem while self-isolated at home. The author also would like to thank Yan Wang at USTC for applying the code of extracting text lines from original SCUT-HCCDoc documents.